# FlowHFT: Flow Policy Induced Optimal High-Frequency Trading under Diverse Market Conditions

### v1

@@ Abstract
High-frequency trading (HFT) is an investing strategy that continuously monitors market states and places bid and ask orders at millisecond speeds. Traditional HFT approaches fit models with historical data and assume that future market states follow similar patterns. This limits the effectiveness of any single model to the specific conditions it was trained for. Additionally, these models achieve optimal solutions only under specific market conditions, such as assumptions about stock price’s stochastic process, stable order flow, and the absence of sudden volatility. Real-world markets, however, are dynamic, diverse, and frequently volatile. To address these challenges, we propose the FlowHFT, a novel imitation learning framework based on flow matching policy. FlowHFT simultaneously learns strategies from numerous expert models, each proficient in particular market scenarios. As a result, our framework can adaptively adjust investment decisions according to the prevailing market state. Furthermore, FlowHFT incorporates a grid-search fine-tuning mechanism. This allows it to refine strategies and achieve superior performance even in complex or extreme market scenarios where expert strategies may be suboptimal. We test FlowHFT in multiple market environments. We first show that flow matching policy is applicable in stochastic market environments, thus enabling FlowHFT to learn trading strategies under different market conditions. Notably, our single framework consistently achieves performance superior to the best expert for each market condition.
@@ 1 Introduction
Market making (MM) is a trading strategy where participants place both buy and sell orders to profit from difference in the bid-ask spread. By providing liquidity through these actions, market makers also enhance market efficiency  madhavan2000market ; harris2002trading . High-frequency trading (HFT) is one type of market making that executes orders at high speeds, often measured in milliseconds or microseconds. This rapid execution allows traders to earn profit on very small price changes within short timeframes.  avellaneda2008high ; cartea2014buy ; ait2017high .
@@
Reinforcement learning (RL) offers another approach to high-frequency trading, framing the problem as an agent (the trading algorithm) interacting with an environment (the market). The agent observes the market state, takes actions and receives rewards based on metrics like profit and loss. Through this interaction, the RL aims to learn a policy—a strategy—that maximizes cumulative rewards. However, RL often focuses on optimizing only single-step actions, which can lead to compounding error  zheng2024reinforcement ; ganesh2018deep  kober2013reinforcement . for instance, slightly suboptimal actions at each step might accumulate, resulting in a large unwanted position. Effective HFT necessitates a long-term view, because profitability typically relies on accumulating numerous small gains from individual steps. Moreover, each action taken interacts with the market environment, thereby impacting the feasibility and profitability of subsequent moves.
@@
To address the challenges associated with traditional HFT methods, we propose FlowHFT, a novel imitation learning framework based on flow matching policy  sohl2015deep ; ho2020denoising ; song2020score  janner2022planning ; wang2022diffusion . This framework consists of two primary components. The first component utilizes imitation learning to develop a pre-trained model for observing the market state and generate corresponding trading actions. We simulate diverse market scenarios and evaluate various traditional models within each. For every scenario, the model achieving the best performance is designated as the ’expert’. FlowHFT then learns by imitating the actions of these designated experts across all scenarios, effectively integrating the strongest known strategy for each market condition into one single, comprehensive model. Imitation learning is useful because these models can indeed generate some optimal solutions in some scenarios. Even in those scenarios where they only generate suboptimal solutions, FlowHFT can still learn their strategies and improve on them subsequently. The second component uses a grid-search mechanism to fine-tune the action proposed by the pre-trained model. The initial action proposed by the pre-trained model, refined by the adjustment from the fine-tuning stage, constitutes the final action.
@@
To the best of our knowledge, this study represents the first application of flow matching policy to a financial stochastic control problem. The results show that the FlowHFT framework integrates knowledge from multiple expert strategies into a single, adaptive model. When market scenarios shift, the framework demonstrates adaptability by generating actions appropriate to the prevailing conditions, effectively leveraging patterns learned from relevant expert demonstrations. Additionally, FlowHFT is designed for rapid inference, enabling the generation of trading actions at millisecond speeds for HFT tasks. Furthermore, the grid search mechanism refine these generated actions. Experimental results indicate that FlowHFT consistently achieves superior performance compared to the best results obtained from baseline models individually optimized for each specific market scenario.
@@
The results indicate that our model exhibits significant robustness to market conditions involving sudden price jumps. We attribute this resilience primarily to our method’s output: it generates sequences of actions over a planning horizon, rather than isolated single-step decisions. By producing action sequences, the model inherently considers near-term trajectory and allows for smoother adjustments, effectively taking into account the broader consequences of initial moves. This sequential planning perspective helps mitigate the compounding error, aligning actions more effectively over time and thus enhancing strategic stability and performance during volatile events like jumps.
@@ 2 Preliminaries
In this section, we formulate the high frequency trading task mathematically. High-frequency trading utilizes automated algorithms for order execution at extremely high speeds, often operating within millisecond or microsecond timescales. The objective is to continuously place buy (bid) and sell (ask) limit orders for a specific financial instrument, while aiming to profit from the bid-ask spread.
@@ 3.1 Backbone of FlowHFT: Flow Matching Policy
Our objective is to train a neural network $v_{\theta}(a,t\mid O_{t})$ to parameterize a conditional vector field. This vector field defines a probability flow that transports samples from a simple prior distribution $p_{\mathrm{prior}}(a_{0})$ (e.g., Gaussian noise) to the distribution of expert trading action sequences $p_{\mathrm{expert}}(a_{E}\mid O_{t})$, conditioned on the observed market state $O_{t}$. Following seminal work on flow matching, we define a time-dependent probability path $p_{t}(a\mid a_{0},a_{E},O_{t})$ that interpolates between an initial random action sequence $a_{0}$ and an expert action sequence $a_{E}$. A simple linear interpolation is given by:
$a_{t}=(1-t)\,a_{0}+t\,a_{E},\quad u_{t}(a_{t}\mid a_{0},a_{E})=a_{E}-a_{0}.$ (2)
The training procedure (Algorithm 1) minimizes the discrepancy between the network’s predicted vector field $v_{\theta}(a_{t},t\mid O_{t})$ and the target vector field $u_{t}(a_{t}\mid a_{0},a_{E})$. Specifically, given a market state $O_{t}$, the model samples a random initial action sequence $a_{0}\sim p_{\mathrm{prior}}$ and a corresponding expert action sequence $a_{E}\sim p_{\mathrm{expert}}(\cdot\mid O_{t})$. It then samples a time $t\sim\mathcal{U}(0,1)$ and forms an interpolated action sequence $a_{t}$ using the defined path. The network is trained by minimizing the following Flow Matching (FM) loss:
@@
At inference (Algorithm 2), to generate an action sequence conditioned on a live market observation $O_{t}$, the policy starts with an action sequence $a_{0}\sim p_{\mathrm{prior}}$. It then iteratively refines this sequence by solving the ordinary differential equation: $\frac{da}{dt}=v_{\theta}(a,t\mid O_{t}),$ from $t=0$ to $t=1$. This is typically achieved using a numerical ODE solver (e.g., Euler or Heun method) over a small number of discretization steps $N$:
@@
This process transforms the initial noise sequence into a coherent, market-aware sequence of high-frequency trading actions.
@@ 8       $t_{\text{current}}\leftarrow t_{\text{current}}+\Delta t_{\text{step}}$;
3.2 Accelerated Inference via Shortcut Policy
While the flow matching policy described in Section 3.1 can accurately model the desired conditional distribution $p(A_{t+1}|O_{t})$, the iterative nature of ODE solving (as in Eq. (4)) with many steps $N$ might be too slow for the stringent low-latency requirements of high-frequency trading (HFT). To address this, we choose shortcut policy frans2024one , denoted $s_{\phi}(a,t,\Delta t|O_{t})$, which is specifically trained to generate high-quality action sequences in significantly fewer discretization steps. This policy is conditioned not only on the state $a$, time $t$, and observation $O_{t}$, but also on the intended integration step size $\Delta t$.
The training of this shortcut policy $s_{\phi}$ (Algorithm 3, where $s_{\theta}$ in the algorithm represents $s_{\phi}$ with parameters $\phi$) is designed to enable it to take larger, more informed discrete steps. The procedure combines two key strategies, leveraging insights from rectified flow and consistency models.
Firstly, for a portion of the training data (Algorithm 3, lines 4-7), the policy is guided by the direct target vector $x_{1}-x_{0}$ (analogous to $a_{E}-a_{0}$ in our action space context). Here, $x_{0}\sim\mathcal{N}(0,I)$ is a noise sample, $x_{1}\sim D$ is an expert sample, and $x_{t}$ is an interpolation. By setting $d=0$ (line 7) for these elements, the model $s_{\phi}(x_{t},t,0|O_{t})$ is implicitly trained to align with this direct path velocity, encouraging straighter, more efficient generation paths.
Secondly, for other batch elements (Algorithm 3, lines 8-12), the policy is trained using a self-consistency objective inspired by methods like progressive distillation. The model $s_{\phi}(x_{t},t,2\Delta t_{\text{small}}|O_{t})$ is trained to predict a velocity whose application over a larger step $2\Delta t_{\text{small}}$ matches the outcome of two consecutive smaller steps, each of size $\Delta t_{\text{small}}$. The target for this larger step is derived from stop-graded predictions of $s_{\phi}$ for these smaller constituent steps: $s_{\text{target}}\leftarrow\text{stopgrad}(s_{t}+s_{t+d})/2$, where $s_{t}$ and $s_{t+d}$ are velocities for the small steps and $d$ here represents $\Delta t_{\text{small}}$. This bootstrapping mechanism enables the policy to learn how to make larger, yet accurate, discrete jumps. The parameter $(d,t)$ in Algorithm 3 (line 2) allows sampling various small step sizes $d$ and interpolation times $t$ for this consistency training.
Once trained, this shortcut policy $s_{\phi}$ can generate action sequences with significantly reduced computational cost using a sampling procedure like Algorithm 4 (where $s_{\theta}$ in the algorithm is the trained $s_{\phi}$). The number of steps $M$ in Algorithm 4 can now be set to a small integer (e.g., 1 to 5), drastically reducing inference latency while aiming to maintain high generation quality due to the specialized training of $s_{\phi}$.
1:  Input: Dataset $D$, model $s_{\theta}$, batch size $B$, parameter $p$, market observations dataset $O$
2:  repeat
3:     Sample batch $\{x_{1}^{i}\}_{i=1}^{B}\sim D$, $\{x_{0}^{i}\}_{i=1}^{B}\sim\mathcal{N}(0,I)$, $\{(d^{i},t^{i})\}_{i=1}^{B}\sim p(d,t)$, $\{O_{t}^{i}\}_{i=1}^{B}\sim O$
4:     for $i=1$ to $\lfloor p\cdot B\rfloor$ do
5:        $x_{t}^{i}\leftarrow(1-t^{i})\cdot x_{0}^{i}+t^{i}\cdot x_{1}^{i}$
6:        $s_{\text{target}}^{i}\leftarrow(x_{1}^{i}-x_{0}^{i})$
7:        $d^{i}\leftarrow 0$
8:     end for
9:     for $i=\lfloor p\cdot B\rfloor+1$ to $B$ do
10:        $x_{t}^{i}\leftarrow(1-t^{i})\cdot x_{0}^{i}+t^{i}\cdot x_{1}^{i}$
11:        $s_{t}^{i}\leftarrow s_{\theta}(x_{t}^{i},t^{i},d^{i}|O_{t}^{i})$Add market observation conditioning}
12:        $x_{t+d}^{i}\leftarrow x_{t}^{i}+d^{i}\cdot s_{t}^{i}$Follow ODE trajectory}
13:        $s_{t+d}^{i}\leftarrow s_{\theta}(x_{t+d}^{i},t^{i}+d^{i},d^{i}|O_{t}^{i})$
14:        $s_{\text{target}}^{i}\leftarrow\text{stopgrad}((s_{t}^{i}+s_{t+d}^{i})/2)$Modified stopgrad position}
15:     end for
16:     Update $\theta$ by minimizing $\frac{1}{B}\sum_{i=1}^{B}\|s_{\theta}(x_{t}^{i},t^{i},d^{i}|O_{t}^{i})-s_{\text{target}}^{i}\|^{2}$Use consistent step size}
17:  until convergence
Algorithm 3 Shortcut Policy Training
1:  Input: Market observation $O_{t}$, trained model $s_{\theta}$, steps $M$
2:  Output: Trading action sequence $a_{M}$
3:  Sample $a_{0}\sim\mathcal{N}(0,I)$
4:  $d\leftarrow 1/M$Keep step size consistent with training}
5:  for $k=0$ to $M-1$ do
6:     $t_{k}\leftarrow k\cdot d$
7:     $a_{k+1}\leftarrow a_{k}+d\cdot s_{\theta}(a_{k},t_{k},d|O_{t})$Use same step size d}
8:  end for
9:  return  $a_{M}$ as the generated trading action sequence
Algorithm 4 Shortcut Policy Inference
@@ 3.3 Fine-tune Pretrained Model
4 Generation of Imitation Learning Data
4.1 Market Simulation and Expert Models
@@
FlowHFT is an imitation learning framework. Therefore, it is crucial to prepare high-quality learning material. In this section, we firstly detail how to generate a set of various market scenarios. Secondly, for each individual scenario, we evaluate a pool of candidate experts—including established traditional algorithms and reinforcement learning agents—and designate the one exhibiting the best performance as the ’expert’ for that context. We use the actions generated by the expert’s strategy to create state-action pairs for training FlowHFT.
@@
We model the mid-price $S_{t}$ as a jump-diffusion process merton1976option , capturing both continuous price movements and sudden jumps. The stochastic differential equation (SDE) governing the mid-price is given by: $dS_{t}=S_{t^{-}}\left(\mu dt+\sigma dB_{H}(t)\right)+S_{t^{-}}(e^{J}-1)dN_{t}$ , where $\mu$ represents the drift coefficient, signifying the expected return of the asset. The asset’s volatility is denoted by $\sigma$, which quantifies the fluctuations in its price. The term $dB_{H}(t)$ corresponds to the increment of a fractional brownian motion process, capturing the continuous component of price movement. Furthermore, $J$ denotes the size of the jump, following a normal distribution $N(\mu_{J},\sigma_{J}^{2})$, which specifies the mean and variance of the jump. Lastly, $dN_{t}$ is a Poisson process with intensity $\lambda_{J}$, modeling the occurrence of jumps, thereby reflecting sudden shifts in the asset’s price.
@@
Next we discuss simulation of order arrival. Order arrivals are modeled using mutually exciting Hawkes processes hawkes1971point ; bacry2015hawkes , effectively capturing the self-exciting and the cross-exciting nature of financial markets. ’self-exciting’ effects indicate that prior buy (or sell) orders increase subsequent arrivals of the same order type, respectively. ’Cross-exciting’ effects describe how buy orders influence sell order arrivals and vice-versa.
@@
The intensity of buy and sell orders, $\lambda_{b}(t)$ and $\lambda_{a}(t)$ respectively, are defined as: $\lambda_{a}(t)=\mu_{a}+\sum_{t_{i}\in\mathcal{N}_{a}}\alpha_{aa}e^{-\beta(t-t_{i})}+\sum_{t_{j}\in\mathcal{N}_{b}}\alpha_{ab}e^{-\beta(t-t_{j})}$ and $\lambda_{b}(t)=\mu_{b}+\sum_{t_{i}\in\mathcal{N}_{b}}\alpha_{bb}e^{-\beta(t-t_{i})}+\sum_{t_{j}\in\mathcal{N}_{a}}\alpha_{ba}e^{-\beta(t-t_{j})}$ where, $\mu_{b}$ and $\mu_{a}$ are the baseline intensities for buy and sell orders, respectively. The parameters $\alpha_{bb}$ and $\alpha_{aa}$ describe the self-exciting effects of buy and buy, and sell and sell orders, respectively, while $\alpha_{ba}$ and $\alpha_{ab}$ represent the cross-exciting effects between buy and sell orders. The parameter $\beta$ denotes the decay rate, which influences how quickly the effects of past events diminish over time. Finally, $N_{b}$ and $N_{a}$ are the sets containing the timestamps of past buy and sell order arrivals, which are used to model the timing of these events.
To generate diverse expert demonstrations for FlowHFT, we utilized a set of expert candidates, including: (1) the Avellaneda-Stoikov (AS) model avellaneda2008high , a foundational stochastic control framework for optimal quoting that considers inventory risk and market volatility; (2) the Guéant-Lehalle-Fernandez-Tapia (GLFT) model gueant2013dealing , which extends such optimal control approaches, often incorporating aspects such as order flow dynamics; (3) a modified GLFT model incorporating a price drift component, designed to adapt strategies in markets exhibiting directional trends; and (4)a model-free Reinforcement Learning (RL) agent trained with Proximal Policy Optimization (PPO) schulman2017proximal , which learns a policy through direct interaction or simulation to maximize a specified reward function. The detailed mathematical formulations, assumptions, and parameterization specifics for each of these expert models are provided in Appendix [Your Appendix Letter/Number].
4.2 Generation of Market Scenarios
@@
We create market scenarios with varying levels of liquidity—categorized as high, medium, and low. Furthermore, to test the model performance under stress, we introduce scenarios featuring sudden market changes and periods of significantly increased volatility, allowing us to examine how effectively each model responds to dynamic and challenging environments. We simulate different market environments by adjusting the following Simulation parameters:
  1. 1.
    Hurst Component Levels: The Hurst component is determined by the Hurst exponent H, which quantifies the long-term memory and self-similar characteristics of the mid-price time series. A large $H>0.5$ indicates trend-following behavior, a small $H<0.5$ indicates mean-reversion, and $H=0.5$ signifies Markov property of memorylessness.
  2. 2.
    Volatility Levels: Controlled by the volatility parameter $\sigma$ in the mid-price dynamics. High volatility corresponds to larger $\sigma$, inducing more significant price fluctuations.
  3. 3.
    Liquidity Levels: Adjusted through the baseline intensities $\mu_{b}$, $\mu_{a}$, and the excitation parameters in the Hawkes processes. High liquidity corresponds to high arrival rates and frequent orders, while low liquidity corresponds to low arrival rates and infrequent orders.
  4. 4.
    Sudden Market Changes: Introduced by incorporating jump events in the mid-price dynamics with higher intensity $\lambda_{J}$ and larger jump sizes $J$.
In the model, the time step sizes are denoted by $dt$ and are within the range 0.01, 0.02. The drifts, represented by $\mu$, vary among 0.01, 0.05, 0.2. The volatilities, $\sigma$, are considered at values 0.1, 0.2, 0.4. The Hurst Exponent, $H$, is evaluated at 0.3, 0.5, 0.7, indicating the level of long-term memory of the time series. The jump intensities, $\lambda_{j}$, are set at 0.05, 0.1, 0.2. Baseline intensities for buying and selling, $\mu_{a}$ and $\mu_{b}$, are equal and chosen from 10, 20, 40. The strategy employed is the "Dominant Stochastic Strategy," which guides the decision-making process within this stochastic framework. Combining these parameters yields $2\times 3\times 3\times 3\times 3\times 3\times 1=486$ parameter combinations.

### v2

# FlowHFT: Imitation Learning via Flow Matching Policy for Optimal High-Frequency Trading under Diverse Market Conditions

@@ Abstract
High-frequency trading (HFT) is an investment strategy that continuously monitors market conditions and places bid and ask orders at nanosecond speeds. Traditional HFT approaches fit models using historical data and assume future market states will follow similar patterns, limiting the effectiveness of any single model to the specific conditions under which it was trained. Additionally, such models achieve optimal performance only under restrictive assumptions, including specific stochastic processes for asset prices, stable order flows, and the absence of sudden volatility spikes. However, real-world financial markets are dynamic, heterogeneous, and frequently volatile. To address these challenges, we propose FlowHFT, a novel imitation learning framework based on a flow matching policy. FlowHFT simultaneously learns strategies from numerous expert models, each of which is proficient in specific market scenarios. As a result, our framework adaptively adjusts investment decisions according to prevailing market conditions. Furthermore, FlowHFT incorporates a grid-search fine-tuning mechanism, enabling it to refine strategies and achieve superior performance even under complex or extreme market conditions where expert strategies may become suboptimal. We evaluate FlowHFT across various market environments. Our study is the first to demonstrate the successful application of flow matching policy within stochastic market environments. Our findings show that FlowHFT can simultaneously learn trading strategies for each market condition in a single model. Notably, our unified framework consistently outperforms the best-performing expert strategy in each tested market scenario.
@@ 1 Introduction
Market making (MM) is a trading strategy where participants place both buy and sell orders to profit from difference in the bid-ask spread. By providing liquidity through these actions, market makers also enhance market efficiency  madhavan2000market ; harris2002trading . High-frequency trading (HFT) is one type of market making that executes orders at high speeds, often measured in milliseconds or microseconds. This rapid execution allows traders to earn profit on very small price changes within short timeframes  avellaneda2008high ; cartea2014buy ; ait2017high .
@@
In the recent literature, Reinforcement learning (RL) has been applied to high-frequency trading, framing the problem as an agent (the trading algorithm) interacting with an environment (the market). The agent observes the market state, takes actions and receives rewards based on metrics like profit and loss. Through this interaction, the RL aims to learn a policy—a strategy—that maximizes cumulative rewards. However, RL often focuses on optimizing only single-step actions, which can lead to compounding error  ganesh2018deep ; kober2013reinforcement ; zheng2024reinforcement for instance, slightly suboptimal actions at each step might accumulate, resulting in a large unwanted position. Effective HFT necessitates a long-term view, because profitability typically relies on accumulating numerous small gains from individual steps. Moreover, each action taken interacts with the market environment, thereby impacting the feasibility and profitability of subsequent moves.
@@
To address the challenges associated with traditional HFT methods, we propose FlowHFT, a novel imitation learning framework based on flow matching policy  chi2024diffusionpolicy ; sohl2015deep ; ho2020denoising ; song2020score ; janner2022planning ; wang2022diffusion . This framework consists of two primary components. The first component utilizes imitation learning to develop a pre-trained model for observing the market state and generate corresponding trading actions. We simulate diverse market scenarios and evaluate various traditional models within each. For every scenario, the model achieving the best performance is designated as the "expert". FlowHFT then learns by imitating the actions of these designated experts across all scenarios, effectively integrating the strongest known strategy for each market condition into one single, comprehensive model. Imitation learning is useful because these models can in fact generate optimal solutions in certain scenarios. Even in those scenarios where they only generate suboptimal solutions, FlowHFT can still learn their strategies and improve on them subsequently. The second component uses a grid search mechanism to fine-tune the action proposed by the pre-trained model. The initial action proposed by the pre-trained model, refined by the adjustment from the fine-tuning stage, constitutes the final action.
@@
To the best of our knowledge, this paper presents the first flow matching policy based imitation learning approach to a financial stochastic control problem. The results show that the FlowHFT framework integrates the knowledge of multiple expert strategies into a single adaptive model. When market scenarios change, the framework demonstrates adaptability by generating actions appropriate to the prevailing conditions, effectively leveraging patterns learned from relevant expert demonstrations. Additionally, FlowHFT is designed for rapid inference, enabling the generation of trading actions at millisecond speeds for HFT tasks. Furthermore, the grid search mechanism refines these generated actions. The experimental results indicate that FlowHFT consistently achieves superior performance compared to the best results obtained from the baseline models individually optimized for each specific market scenario.
@@
The results indicate that our model exhibits significant robustness to market conditions involving sudden price jumps. We attribute this resilience primarily to our method’s output: it generates sequences of actions over a planning horizon, rather than isolated single-step decisions. By producing action sequences, the model inherently considers a near-term trajectory and allows for smoother adjustments, effectively taking into account the broader consequences of initial moves. This sequential planning perspective helps mitigate the compounding error, aligning actions more effectively over time and thus enhancing strategic stability and performance during volatile events like jumps.
Figure 1: Visualization of FlowHFT’s Input Observation Sequence and Output Action Sequence
@@ 2 Preliminaries
In this section, we formulate the high frequency trading task mathematically. High-frequency trading utilizes automated algorithms for order execution at extremely high speeds, often operating within millisecond or nanosecond timescales. The objective is to continuously place buy (bid) and sell (ask) limit orders for a specific financial instrument, while aiming to profit from the bid-ask spread.
@@ 3.1 Backbone of FlowHFT: Flow Matching Policy
Figure 2: This figure illustrates the training of a flow matching policy. The inputs are various market situations. For each situation, the best expert is chosen from a pool of models. These optimal market-expert pairings are then used to train the flow matching policy, which operates as a vector field. This field maps the distribution of market conditions to a corresponding optimal action distribution.
@@
Our objective is to train a neural network $v_{\theta}(a,t\mid O_{t})$ to parameterize a conditional vector field. This vector field defines a probability flow that transports samples from a simple prior distribution $p_{\mathrm{prior}}(a_{0})$ (e.g., Gaussian noise) to the distribution of expert trading action sequences $p_{\mathrm{expert}}(a_{E}\mid O_{t})$, conditioned on the observed market state $O_{t}$. Following seminal work on flow matching, we define a time-dependent probability path $p_{t}(a\mid a_{0},a_{E},O_{t})$ that interpolates between an initial random action sequence $a_{0}$ and an expert action sequence $a_{E}$. A simple linear interpolation is given by: $a_{t}=(1-t)\,a_{0}+t\,a_{E},\quad u_{t}(a_{t}\mid a_{0},a_{E})=a_{E}-a_{0}$. The training procedure (Algorithm 1) minimizes the discrepancy between the network’s predicted vector field $v_{\theta}(a_{t},t\mid O_{t})$ and the target vector field $u_{t}(a_{t}\mid a_{0},a_{E})$. Specifically, given a market state $O_{t}$, the model samples a random initial action sequence $a_{0}\sim p_{\mathrm{prior}}$ and a corresponding expert action sequence $a_{E}\sim p_{\mathrm{expert}}(\cdot\mid O_{t})$. It then samples a time $t\sim\mathcal{U}(0,1)$ and forms an interpolated action sequence $a_{t}$ using the defined path. The network is trained by minimizing the following Flow Matching (FM) loss:
@@ 12      $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$;
At inference (Algorithm 2), to generate an action sequence conditioned on a live market observation $O_{t}$, the policy starts with an action sequence $a_{0}\sim p_{\mathrm{prior}}$. It then iteratively refines this sequence by solving the ordinary differential equation: $\frac{da}{dt}=v_{\theta}(a,t\mid O_{t}),$ from $t=0$ to $t=1$. This is typically achieved using a numerical ODE solver (e.g., Euler or Heun method) over a small number of discretization steps $N$. This process transforms the initial noise sequence into a coherent, market-aware sequence of high-frequency trading actions.
@@ 3.2 Fine-tune Pretrained Model
Figure 3: This figure shows the action generation. The neural network of flow matching policy (left) provides initial bid/ask outputs. A grid search (right) identifies adjustments ($\Delta$bid,$\Delta$ask) to refine these actions. Combine the initial result and adjustment together to get the final results.
4 Data of Imitation Learning
@@
FlowHFT is an imitation learning framework. Therefore, it is crucial to prepare high-quality learning material. In this section, we firstly detail how to generate a set of various market scenarios. Secondly, for each individual scenario, we evaluate a pool of candidate experts—including established traditional algorithms and reinforcement learning agents—and designate the one exhibiting the best performance as the "expert" for that context. We use the actions generated by the expert’s strategy to create state-action pairs for training FlowHFT.
4.1 Market Simulation
We model the mid-price $S_{t}$ as a jump-diffusion process merton1976option , capturing both continuous price movements and sudden jumps. The stochastic differential equation (SDE) governing the mid-price is given by:
$dS_{t}=S_{t^{-}}\left(\mu dt+\sigma dB_{H}(t)\right)+S_{t^{-}}(e^{J}-1)dN_{t}$ (4)
@@ 4.1 Market Simulation
where $\mu$ represents the drift coefficient, signifying the expected return of the asset. The asset’s volatility is denoted by $\sigma$, which quantifies the fluctuations in its price. The term $dB_{H}(t)$ corresponds to the increment of a fractional brownian motion process, capturing the continuous component of price movement. Furthermore, $J$ denotes the size of the jump, following a normal distribution $N(\mu_{J},\sigma_{J}^{2})$, which specifies the mean and variance of the jump. Lastly, $dN_{t}$ is a Poisson process with intensity $\lambda_{J}$, modeling the occurrence of jumps, thereby reflecting sudden shifts in the asset’s price.
@@
Next we discuss simulation of order arrival. Order arrivals are modeled using mutually exciting Hawkes processes bacry2015hawkes ; hawkes1971point , effectively capturing the self-exciting and the cross-exciting nature of financial markets. "self-exciting" effects indicate that prior buy (or sell) orders increase subsequent arrivals of the same order type, respectively. "Cross-exciting" effects describe how buy orders influence sell order arrivals and vice-versa.
The intensity of buy and sell orders, $\lambda_{b}(t)$ and $\lambda_{a}(t)$ respectively, are defined as:
$\lambda_{b}(t)$ $=\mu_{b}+\sum_{t_{i}\in\mathcal{N}_{b}}\alpha_{bb}e^{-\beta(t-t_{i})}+\sum_{t_{j}\in\mathcal{N}_{a}}\alpha_{ba}e^{-\beta(t-t_{j})},$ (5)
$\lambda_{a}(t)$ $=\mu_{a}+\sum_{t_{i}\in\mathcal{N}_{a}}\alpha_{aa}e^{-\beta(t-t_{i})}+\sum_{t_{j}\in\mathcal{N}_{b}}\alpha_{ab}e^{-\beta(t-t_{j})},$ (6)
@@
where, $\mu_{b}$ and $\mu_{a}$ are the baseline intensities for buy and sell orders, respectively. The parameters $\alpha_{bb}$ and $\alpha_{aa}$ describe the self-exciting effects of buy and buy, and sell and sell orders, respectively, while $\alpha_{ba}$ and $\alpha_{ab}$ represent the cross-exciting effects between buy and sell orders. The parameter $\beta$ denotes the decay rate, which influences how quickly the effects of past events diminish over time. Finally, $N_{b}$ and $N_{a}$ are the sets containing the timestamps of past buy and sell order arrivals, which are used to model the timing of these events.
@@
We create market scenarios with varying levels of liquidity—categorized as high, medium, and low. Furthermore, to test the model performance under stress, we introduce scenarios featuring sudden market changes and periods of significantly increased volatility, allowing us to examine how effectively each model responds to dynamic and challenging environments. We simulate different market environments by adjusting the following Simulation parameters. Combining these parameters yields $3\times 3\times 3\times 2\times 2\times 3\times 1=324$ parameter combinations.
Figure 4: Parameters settings of market simulation